Among two supplementry angles, the measure of the large angle 36 degree more than the smaller angle.find their measures.
Answer by nerdybill(6959)

Among two supplementry angles, the measure of the large angle 36 degree more than the smaller angle.find their measures.
.
Let x = measure of smaller angle
then
x+36 = measure of larger angle
.
x + x+36 = 180
2x+36 = 180
2x = 144
x = 72 degrees (smaller angle)
.
Larger angle:
x+36 = 72+36 = 108 degrees (larger angle)